500 mL of 0.0100 M CaCl2 is mixed with 500 mL of 0.0100 M Na2SO4. How many grams of CaSO4 will precipitate when the equilibrium is reached? The solubility product constant of CaSO4 is Ksp = 2.50 × 10-5MM(CaSO4) = 136.14 g mol–1.

  •  A. 1.36 g
  •  B. 2.04 g
  •  C. No precipitate will form
  •  D. 0.68 g
  •  E. 5.45 g
1
Expert's answer
2021-03-17T08:45:51-0400

CaCl2 + Na2SO4 = 2NaCl + CaSO4

СM = n/V

n = CM x V

n = m/M

n (CaCl2) = 0.0100 x 0.5 = 0.005 mol

n (Na2SO4) = 0.0100 x 0.5 = 0.005 mol

According to the equation, n (CaCl2) = n (Na2SO4) = n (CaSO4) = 0.005 mol

M (CaSO4) = 136.14 g/mol

m (CaSO4) = n x M = 0.005 x 136.14 = 0.68 g


Answer: D - 0.68 g